X=2 -5x, x(x) = -2 z LI -2 5
Answer #1

Ques: solve the System -5 |x, X (T) = [-2] -2 Ansi- The Given System X (1) of the form A= Now So 2 -5 where (? - ] we find tdi= Now & U2=1 so Eigenvalue of Ais and daai find Eigenvector Corresponding to hizi (A - l.I) u=0 { 13:21:15:) }6) - 609 ( 2-(21) Now Eigen vector Corresponding to do = -i, U- Now Solution X(t)= G M eart X(t) = G( 2 che + Gp4te trisint) +G to ve = [C .-2i = 12+5 i 12+5 i -Qi Xi G 5 12 +12i -) 6= -5 +126 co 2 Put value of Ca in (uis weget -G+/- ( -5 +126)] = 5 5 + -5+12i 2. tg ug e dat - t Co Une ant Hence solution of the system 2 Cost +29 Cost -5 Cost + 12 Cost is the required solution Note :-